OpenStudy (anonymous):

Solve (x/3) + 2 > 8.

OpenStudy (anonymous):

\[\frac{x}{3} + 2 > 8\] First, subtract 2 from both sides. \[\frac{x}{3}+ 2 (-2) > 8 - 2\] \[\frac{x}{3} > 6\] Then, multiply both sides by 3. \[\frac{x}{3(3)} > 6(3)\] \[x = 18\]

OpenStudy (anonymous):

thank you :)